# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v1

@@ Abstract
We extend the fourth order, two stage Multi-Derivative Runge Kutta (MDRK) scheme of Li and Du to the Flux Reconstruction (FR) framework by writing both of the stages in terms of a time averaged flux and then use the approximate Lax-Wendroff procedure. Numerical flux is computed in each stage using D2 dissipation and EA flux, enhancing Fourier CFL stability and accuracy respectively. A subcell based blending limiter is developed for the MDRK scheme which ensures that the limited scheme is provably admissibility preserving. Along with being admissibility preserving, the blending scheme is constructed to minimize dissipation errors by using Gauss-Legendre solution points and performing MUSCL-Hancock reconstruction on subcells. The accuracy enhancement of the blending scheme is numerically verified on compressible Euler’s equations, with test cases involving shocks and small-scale structures.
@@ 1 Introduction
The idea to apply Multiderivative Runge-Kutta Flux Reconstruction on the time dependent conservation law
$\bm{u}_{t}+\bm{f}\left(\bm{u}\right)_{x}=\bm{0}$
$\begin{split}\bm{u}^{\ast}&=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}^{\ast}\\ \bm{u}^{n+1}&=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}\end{split}$
where
$\partial_{x}\bm{F}^{\ast}=\partial_{x}\bm{F}^{\ast}\left(\bm{u}^{n}\right)\approx\partial_{x}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\text{\text{{{d}}}}t,\qquad\partial_{x}\bm{F}=\partial_{x}\bm{F}\left(\bm{u}^{n},\bm{u}^{\ast}\right)\approx\partial_{x}\int_{t^{n}}^{t^{n+1}}\bm{f}\text{\text{{{d}}}}t$
[…]
1.1 Motivation and literature review
Higher order methods incorporate higher order terms and perform more computations per degree of freedom, increasing accuracy and arithmetic intensity, and making these methods relevant to the current state of memory-bound HPC hardware [1, 31]. Spectral element methods like Discontinuous Galerkin (DG) and Flux Reconstruction (FR) are two high order methods that are suitable for modern requirements due to their parallelization, capability of handling curved meshes and shock capturing.
Discontinuous Galerkin (DG) is a Spectral Element Method first introduced by Reed and Hill [24] for neutron transport equations and developed for fluid dynamics equations by Cockburn and Shu and others, see [9] and the references therein. The DG method uses an approximate solution which is a polynomial within each element and is allowed to be discontinuous across interfaces. The neighbouring DG elements are coupled only through the numerical flux and thus bulk of computations are local to the element, minimizing data transfers.
@@ 1.1 Motivation and literature review
Lax-Wendroff schemes perform a Taylor’s expansion in time to the order of the desired accuracy and compute the temporal derivatives locally. Multiderivative Runge-Kutta (MDRK) scheme also make use of temporal derivatives but they combine them with multiple stages to obtain the desired order of accuracy. As MDRK schemes use both temporal derivatives and multiple stages, they are a generalization of LW and standard multistage RK methods [26]. MDRK methods typically require fewer temporal derivatives in contrast to the Lax-Wendroff schemes and fewer stages in contrast to the standard RK methods, which is what makes them promising.
Multiderivative Runge-Kutta (MDRK) [19] methods were initially developed to solve systems of ODE like
@@
that use temporal derivatives of $\bm{L}$, see [26] for a review of MDRK methods for solving ODEs. They were first used for temporal discretization of hyperbolic conservation laws in [26] by using Weighted Essentially Non-Oscillatary (WENO) [29] and Discontinuous Galerkin (DG) [9] methods for spatial discretization.
@@
In [18], a two stage fourth order MDRK scheme was introduced for solving hyperbolic conservation laws by solving a Generalized Riemann Problem (GRP). In this work, we show the first combination of MDRK with a Flux Reconstruction scheme by using the scheme of [18]. We also use the construction of the numerical flux from [3]. In particular, we use the D2 dissipation introduced in [3] and show that it leads to enhanced Fourier CFL stability limiter. We also use the EA scheme from [3] which leads to enhanced accuracy for non-linear problems when using Gauss-Legendre solution points. We also develop admissibility preserving subcell based blending scheme based on [4] and show how it superior to other schemes like a TVB limiter.
1.2 Outline
The rest of this paper is organized as follows. The discretization of the domain and function approximation by polynomials is presented in Section 2. Section 3 reviews the one dimensional Runge-Kutta Flux Reconstruction (RKFR) method and Section 4 introduces the MDRK method in an FR framework. In particular, Section 4.3 discusses approximate Lax-Wendroff procedure applied to MDRK, Sections 4.4 discuss the D2 dissipation for computing the dissipative part of the numerical flux to enhance Fourier CFL stability limit and Section 4.6 discusses the EA scheme for computing the central part of numerical flux to enhance stability. The Fourier stability analysis is performed in Section 5 to demonstrate the improved stability of D2 dissipation. In Section 6, we explain the admissibility preserving blending limiter for MDRK scheme. The numerical results validating the order of accuracy and capability of the blending scheme are shown in Section 7 and a summary of the new MDRK scheme is presented in Section 8.
2 conservation law
@@ 1.2 Outline
where each admissibility constraint $p_{k}$ is concave if $p_{j}>0$ for all $j<k$. For Euler’s equations, $K=2$ and $p_{1},p_{2}$ are density, pressure functions respectively; if the density is positive then pressure is a concave function of the conserved variables. We will divide the computational domain $\Omega$ into disjoint elements $\Omega_{e}$, with
@@
In this work, we will only be using degree $N=3$ polynomials as the Multiderivative Runge-Kutta discretization that we use [18] is fourth order. However, the description is given for general $N$ for ease of notation.
@@ 3 Runge-Kutta FR
The RKFR scheme is based on an FR spatial discretization leading to a system of ODE followed by application of an RK scheme to march forward in time. The key idea is to construct a continuous polynomial approximation of the flux which is then used in a collocation scheme to update the nodal solution values. At some time $t$, we have the piecewise polynomial solution defined inside each cell; the FR scheme can be described by the following steps.
@@
$\bm{f}_{e+\frac{1}{2}}(t)=\bm{f}(\bm{f}_{h}(x_{e+\frac{1}{2}}^{-},t),\bm{f}_{h}(x_{e+\frac{1}{2}}^{+},t))$
@@
$g_{L}(0)=1,\qquad$ $g_{R}(0)=0$
$g_{L}(1)=0,\qquad$ $g_{R}(1)=1$
which ensures the continuity of the flux, i.e., $\bm{f}_{h}(x_{e+\frac{1}{2}}^{-},t)=\bm{f}_{h}(x_{e+\frac{1}{2}}^{+},t)=\bm{f}_{e+\frac{1}{2}}(t)$. The reader is referred to [3, 16, 36] for discussion and choices of correction functions.
@@ 4 Multi-derivative Runge-Kutta FR scheme
The Multi-Derivative Runge-Kutta scheme of [18] will be cast in a Flux Reconstruction framework by writing each stage in terms of a time averaged flux that will be computed with the Approximate Lax-Wendroff procedure [42]. The scheme to solve (1) is thus given by the following two stage procedure
$\bm{u}^{\ast}$ $=\bm{u}^{n}-\frac{\Delta t}{2}\partial_{x}\bm{F}^{\ast}$ (4)
$\bm{u}^{n+1}$ $=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}$ (5)
@@
$\begin{split}\bm{F}^{\ast}:=&\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)\\ \bm{F}:=&\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)+2\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{\ast}\right)\right)=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\bm{f}\left(\bm{u}\right)\text{\text{{{d}}}}t+O(\Delta t^{4})\end{split}$ (6)
The approximations in (6) ensure that we will get the formal fourth order of accuracy. Then the solution update at the nodes is written as a collocation scheme
$\begin{split}\bm{u}_{e,p}^{n+1}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\frac{\text{\text{{{d}}}}\bm{F}^{\ast}_{h}}{\text{\text{{{d}}}}\xi}(\xi_{p})\\ \bm{u}_{e,p}^{n+1}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\frac{\text{\text{{{d}}}}\bm{F}_{h}}{\text{\text{{{d}}}}\xi}(\xi_{p})\end{split},\qquad 0\leq p\leq N$ (7)
where we take $N=3$ to get fourth order accuracy. The major work in the above scheme is involved in the construction of the time average flux approximation $\bm{F}_{h}$ which is explained in subsequent sections.
@@ 4.1 Conservation property
$\begin{split}\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{\ast}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}^{\ast}_{h}}{\partial\xi}(\xi_{p})\\ \sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n+1}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}_{h}}{\partial\xi}(\xi_{p})\end{split}$
The correction functions are of degree $N+1$ and thus the fluxes $\bm{F}^{\ast}_{h},\bm{F}_{h}$ are polynomials of degree $\leq N+1$. If the quadrature is exact for polynomials of degree atleast $N$, which is true for both GLL and GL points, then the quadrature is exact for the flux derivative term and we can write it as an integral, which leads to
$\begin{split}\int_{\Omega_{e}}\bm{u}^{\ast}_{h}\text{\text{{{d}}}}x&=\int_{\Omega_{e}}\bm{u}_{h}^{n}\text{\text{{{d}}}}x-\Delta t[\bm{F}^{\ast}_{e+\frac{1}{2}}-\bm{F}^{\ast}_{e-\frac{1}{2}}]\\ \int_{\Omega_{e}}\bm{u}^{n+1}_{h}\text{\text{{{d}}}}x&=\int_{\Omega_{e}}\bm{u}_{h}^{n}\text{\text{{{d}}}}x-\Delta t[\bm{F}_{e+\frac{1}{2}}-\bm{F}_{e-\frac{1}{2}}]\end{split}$ (8)
@@
$\bm{u}_{e,p}^{n}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}\quad\forall e,j\qquad\Longrightarrow\qquad\bm{u}_{e,p}^{n+1}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}\quad\forall e,p$
@@ 4.2 Reconstruction of the time average flux
To complete the description of the MDRK method (7), we must explain the method for the computation of the time average fluxes $\bm{F}^{\ast}_{h},\bm{F}_{h}$ when evolving from $t^{n}$ to $t^{n+1}$. In the first stage (4), we compute $\bm{F}^{\ast}_{h}$ which is then used to evolve to $\bm{u}^{\ast}$. In the second stage (5), $\bm{u}^{n},\bm{u}^{\ast}$ are used to compute $\bm{F}_{h}$ which is used from evolution to $\bm{u}^{n+1}$. Since the reconstruction procedure is same for both of the time average fluxes, we explain the procedure of Flux Reconstruction of $\bm{F}_{h}(\xi)$, which is a time averaged flux over the interval $[t^{n},t^{n+1}]$, performed in three steps. Note that all results in this work use degree $N=3$, although the following steps are written for general $N$.
@@
Use the approximate Lax-Wendroff procedure of [42] to compute the time average flux $\bm{F}$ at all the solution points
$\bm{F}_{p}^{e}\approx\bm{F}(\xi_{p}),\qquad 0\leq p\leq N$
The approximate LW procedure is the subsection Section 4.3.
@@
$\bm{F}_{h}^{\delta}(\xi)=\sum_{j=0}^{N}\bm{F}_{j}^{e}\ell_{p}(\xi)$
which however may not be continuous across the elements. This is illustrated in Figure (1b).
@@
Modify the flux approximation $\bm{F}_{h}^{\delta}(\xi)$ so that it becomes continuous across the elements. Let $\bm{F}_{e+\frac{1}{2}}$ be some numerical flux function that approximates the flux $\bm{F}$ at $x=x_{e+\frac{1}{2}}$. Then the continuous flux approximation is given by
$\bm{F}_{h}(\xi)=\left[\bm{F}_{e-\frac{1}{2}}-\bm{F}_{h}^{\delta}(0)\right]g_{L}(\xi)+\bm{F}_{h}^{\delta}(\xi)+\left[\bm{F}_{e+\frac{1}{2}}-\bm{F}_{h}^{\delta}(1)\right]g_{R}(\xi)$
@@
$\mathsf{\partial_{\xi}F_{h}}=\left[F_{e}-\frac{1}{2}-\mathsf{V}_{L}^{\top}\mathsf{F}\right]\mathsf{b}_{L}+\mathsf{D}\mathsf{F}+\left[F_{e}+\frac{1}{2}-\mathsf{V}_{R}^{\top}\mathsf{F}\right]\mathsf{b}_{R},\qquad\mathsf{b}_{L}=\begin{bmatrix}g_{L}^{\prime}(\xi_{0})\\ \vdots\\ g_{L}^{\prime}(\xi_{N})\end{bmatrix},\qquad\mathsf{b}_{R}=\begin{bmatrix}g_{R}^{\prime}(\xi_{0})\\ \vdots\\ g_{R}^{\prime}(\xi_{N})\end{bmatrix}$
@@
$\mathsf{\partial_{\xi}F_{h}}=\bm{F}_{e-\frac{1}{2}}\mathsf{b}_{L}+\mathsf{D}_{1}\mathsf{F}+\bm{F}_{e+\frac{1}{2}}\mathsf{b}_{R},\qquad\mathsf{D}_{1}=\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{\top}-\mathsf{b}_{R}\mathsf{V}_{R}^{\top}$ (10)

### v2

@@ Abstract
We extend the fourth order, two stage Multi-Derivative Runge Kutta (MDRK) scheme to the Flux Reconstruction (FR) framework by writing both stages in terms of a time averaged flux and then using the approximate Lax-Wendroff procedure to compute the time averaged flux. Numerical flux is carefully constructed to enhance Fourier CFL stability and accuracy. A subcell based blending limiter is developed for the MDRK scheme which ensures that the limited scheme is provably admissibility preserving. Along with being admissibility preserving, the blending scheme is constructed to minimize dissipation errors by using Gauss-Legendre solution points and performing MUSCL-Hancock reconstruction on subcells. The accuracy enhancement of the blending scheme is numerically verified on compressible Euler’s equations, with test cases involving shocks and small-scale structures.
[…]
@@ 1 Introduction
Higher order methods incorporate higher order terms and perform more computations per degree of freedom, increasing accuracy and arithmetic intensity, and making these methods relevant to the current state of memory-bound HPC hardware [1, 31]. Spectral element methods like Discontinuous Galerkin (DG) and Flux Reconstruction (FR) are two high order methods that are suitable for modern requirements due to their parallelization, capability of handling curved meshes and shock capturing. Discontinuous Galerkin (DG) is a Spectral Element Method first introduced by Reed and Hill [24] for neutron transport equations and developed for fluid dynamics equations by Cockburn and Shu and others, see [9] and the references therein. The DG method uses an approximate solution which is a polynomial within each element and is allowed to be discontinuous across element interfaces. The neighbouring DG elements are coupled only through the numerical flux and thus bulk of computations are local to the element, minimizing data transfers.
@@
Lax-Wendroff schemes perform a Taylor’s expansion of the solution in time to the order of the desired accuracy and compute the temporal derivatives locally. Multiderivative Runge-Kutta (MDRK) schemes also use temporal derivatives but combine them with multiple stages to obtain the desired order of accuracy. As MDRK schemes use both temporal derivatives and multiple stages, they are a generalization of LW and standard multistage RK methods [26]. MDRK methods typically require fewer temporal derivatives in contrast to the Lax-Wendroff schemes and fewer stages in contrast to the standard RK methods, which is what makes them promising.
For an ODE of the form
@@
which is obtained after spatial discretization of a PDE, Runge-Kutta methods make use of only the right hand side function $\bm{L}$. Multiderivative Runge-Kutta (MDRK) [19] methods were initially developed to solve systems of ODE that also make use temporal derivatives of $\bm{L}$, see [26] for a review of MDRK methods for solving ODEs. They were first used for temporal discretization of hyperbolic conservation laws in [26] by using Weighted Essentially Non-Oscillatary (WENO) [29] and Discontinuous Galerkin (DG) [9] methods for spatial discretization.
@@
In [18], a two stage fourth order MDRK scheme was introduced for solving hyperbolic conservation laws by solving a Generalized Riemann Problem (GRP). In this work, we show the first combination of MDRK with a Flux Reconstruction scheme by using the scheme of [18]. The idea to apply Multiderivative Runge-Kutta Flux Reconstruction on the time dependent conservation law
$\bm{u}_{t}+\bm{f}\left(\bm{u}\right)_{x}=\bm{0}$
$\begin{split}\bm{u}^{\ast}&=\bm{u}^{n}-\frac{\Delta t}{2}\partial_{x}\bm{F}\\ \bm{u}^{n+1}&=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}^{\ast}\end{split}$
where
$\bm{F}=\bm{F}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\text{\text{{{d}}}}t,\qquad\bm{F}^{\ast}=\bm{F}^{\ast}\left(\bm{u}^{n},\bm{u}^{\ast}\right)\approx\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\bm{f}\text{\text{{{d}}}}t$
The method is two-stage; in the first stage, $\bm{F}$ is locally approximated and then flux reconstruction [16] is used to construct a globally continuous approximation of $\bm{F}$ which is used to perform evolution to $\bm{u}^{\ast}$; and the same procedure is then performed using $\bm{F}^{\ast}$ for evolution to $\bm{u}^{n+1}$. We also use the construction of the numerical flux from [3]; in particular, we use the D2 dissipation introduced in [3] and show that it leads to enhanced Fourier CFL stability limit. We also use the EA scheme from [3] which leads to enhanced accuracy for non-linear problems when using Gauss-Legendre solution points. We also develop admissibility preserving subcell based blending scheme based on [4] and show how it superior to other schemes like a TVB limiter.
The rest of this paper is organized as follows. The discretization of the domain and function approximation by polynomials is presented in Section 2. Section 3 reviews the one dimensional Runge-Kutta Flux Reconstruction (RKFR) method and Section 4 introduces the MDRK method in an FR framework. In particular, Section 4.3 discusses approximate Lax-Wendroff procedure applied to MDRK, Section 4.4 discusses the D2 dissipation for computing the dissipative part of the numerical flux to enhance Fourier CFL stability limit and Sections 4.54.6 discuss the EA scheme for computing the central part of numerical flux to enhance stability. The Fourier stability analysis is performed in Section 5 to demonstrate the improved stability of D2 dissipation. In Section 6, we explain the admissibility preserving blending limiter for MDRK scheme. The numerical results validating the order of accuracy and capability of the blending scheme are shown in Section 7 and a summary of the new MDRK scheme is presented in Section 8.
2 Conservation law and solution space
@@
where each admissibility constraint $p_{k}$ is a concave function if $p_{j}>0$ for all $j<k$. For Euler’s equations, $K=2$ and $p_{1},p_{2}$ are density, pressure functions respectively; if the density is positive then pressure is a concave function of the conserved variables.
We will divide the computational domain $\Omega$ into disjoint elements $\Omega_{e}$, with
@@
In this work, we will only be using degree $N=3$ polynomials as the Multiderivative Runge-Kutta discretization that we use [18] is fourth order accurate in time. However, the description is given for general $N$ for ease of notation.
@@ 3 Runge-Kutta FR
We first give a brief resume of RKFR scheme since our method has some similarities and borrows ideas from FR method. The RKFR scheme is based on an FR spatial discretization leading to a system of ODE followed by application of an RK scheme to march forward in time. The key idea is to construct a continuous polynomial approximation of the flux which is then used in a collocation scheme to update the nodal solution values. At some time $t$, we have the piecewise polynomial solution defined inside each cell; the FR scheme can be described by the following steps.
@@
$\bm{f}_{{e+\frac{1}{2}}}(t)=\bm{f}(\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{-},t),\bm{u}_{h}(x_{{e+\frac{1}{2}}}^{+},t))$
@@
$g_{L}(0)=g_{R}(1)=1,\qquad$ $g_{R}(0)=g_{L}(1)=0$
which ensures the continuity of the flux, i.e., $\bm{f}_{h}(x_{{e+\frac{1}{2}}}^{-},t)=\bm{f}_{h}(x_{{e+\frac{1}{2}}}^{+},t)=\bm{f}_{{e+\frac{1}{2}}}(t)$. The reader is referred to [3, 16, 36] for detailed discussion and choices of correction functions.
@@ 4 Multi-derivative Runge-Kutta FR scheme
Multiderivative Runge-Kutta [19] methods were initially developed to solve systems of ODE like
$\frac{\mathrm{d}\bm{u}}{\mathrm{d}t}=\bm{L}(\bm{u})$ (4)
that use temporal derivatives of $\bm{L}$. In this work, we use the two stage fourth order multiderivative Runge-Kutta method from [18]. For the system of ODE (4), the MDRK scheme of [18] to evolve from $t^{n}$ to $t^{n+1}$ is given by
$\begin{split}\bm{u}^{\ast}&=\bm{u}^{n}+\frac{1}{2}\Delta t\bm{L}(\bm{u}^{n})+\frac{\Delta t^{2}}{8}\frac{\mathrm{d}\bm{L}}{\mathrm{d}t}(\bm{u}^{n})\\ \bm{u}^{n+1}&=\bm{u}^{n}+\Delta t\bm{L}(\bm{u}^{n})+\frac{\Delta t^{2}}{6}\left(\frac{\mathrm{d}\bm{L}}{\mathrm{d}t}(\bm{u}^{n})+2\frac{\mathrm{d}\bm{L}}{\mathrm{d}t}(\bm{u}^{\ast})\right)\end{split}$
In order to solve the 1-D conservation law (1) using the above scheme, we formally set $\bm{L}=-\bm{f}(\bm{u})_{x}$ to get the following two stage procedure
$\bm{u}^{\ast}$ $=\bm{u}^{n}-\frac{\Delta t}{2}\partial_{x}\bm{F}$ (5)
$\bm{u}^{n+1}$ $=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}^{\ast}$ (6)
@@
$\begin{split}\bm{F}&:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)\\ \bm{F}^{\ast}&:=\bm{f}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\bm{f}(\bm{u}^{n})+2\frac{\partial}{\partial t}\bm{f}(\bm{u}^{\ast})\right)\end{split}$ (7)
The formal order of accuracy of the scheme (Appendix A) is obtained from
$\partial_{x}\bm{F}^{\ast}=\frac{1}{\Delta t}\partial_{x}\int_{t^{n}}^{t^{n+1}}\bm{f}+O(\Delta t^{4})$
The idea is to use (56) to obtain solution update at the nodes written as a collocation scheme
$\begin{split}\bm{u}^{\ast}_{e,p}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{2\Delta x_{e}}\frac{\text{\text{{{d}}}}\bm{F}_{h}}{\text{\text{{{d}}}}\xi}(\xi_{p})\\ \bm{u}_{e,p}^{n+1}&=\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\frac{\text{\text{{{d}}}}\bm{F}^{\ast}_{h}}{\text{\text{{{d}}}}\xi}(\xi_{p})\end{split},\qquad 0\leq p\leq N$ (8)
where we take $N=3$ to get fourth order accuracy. The major work is in the construction of the time average flux approximations $\bm{F}_{h},\bm{F}^{\ast}_{h}$ which is explained in subsequent sections.
@@ 4.1 Conservation property
$\begin{split}\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{\ast}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{2\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}_{h}}{\partial\xi}(\xi_{p})\\ \sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n+1}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}^{\ast}_{h}}{\partial\xi}(\xi_{p})\end{split}$
The correction functions are of degree $N+1$ and thus the fluxes $\bm{F}_{h},\bm{F}^{\ast}_{h}$ are polynomials of degree $\leq N+1$. If the quadrature is exact for polynomials of degree at least $N$, which is true for both GLL and GL points, then the quadrature is exact for the flux derivative term and we can write it as an integral, which leads to
$\begin{split}\int_{\Omega_{e}}\bm{u}^{\ast}_{h}\text{\text{{{d}}}}x&=\int_{\Omega_{e}}\bm{u}_{h}^{n}\text{\text{{{d}}}}x-\frac{\Delta t}{2}[\bm{F}_{{e+\frac{1}{2}}}-\bm{F}_{{e-\frac{1}{2}}}]\\ \int_{\Omega_{e}}\bm{u}^{n+1}_{h}\text{\text{{{d}}}}x&=\int_{\Omega_{e}}\bm{u}_{h}^{n}\text{\text{{{d}}}}x-\Delta t[\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast}_{{e-\frac{1}{2}}}]\end{split}$ (9)
@@
$\bm{u}_{e,p}^{n}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}\quad\forall e,p\qquad\Longrightarrow\qquad\bm{u}_{e,p}^{n+1}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}\quad\forall e,p$
@@ 4.2 Reconstruction of the time average flux
To complete the description of the MDRK method (8), we must explain the method for the computation of the time average fluxes $\bm{F}_{h},\bm{F}^{\ast}_{h}$ when evolving from $t^{n}$ to $t^{n+1}$. In the first stage (5), we compute $\bm{F}_{h}$ which is then used to evolve to $\bm{u}^{\ast}$. In the second stage (6), $\bm{u}^{n},\bm{u}^{\ast}$ are used to compute $\bm{F}^{\ast}_{h}$ which is used from evolution to $\bm{u}^{n+1}$. Since the reconstruction procedure is same for both of the time average fluxes, we explain the procedure of Flux Reconstruction of $\bm{F}^{\ast}_{h}(\xi)$, which is a time averaged flux over the interval $[t^{n},t^{n+1}]$, performed in three steps. Note that all results in this work use degree $N=3$, although the following steps are written for general $N$.
@@
Use the approximate Lax-Wendroff procedure of [43] to compute the time average flux $\bm{F}^{\ast}$ at all the solution points
$\bm{F}^{\ast}_{e,p}\approx\bm{F}^{\ast}(\xi_{p}),\qquad 0\leq p\leq N$
The approximate LW procedure explained in Section 4.3.
@@
$\bm{F}^{\ast\delta}_{h}(\xi)=\sum_{p=0}^{N}\bm{F}^{\ast}_{e,p}\ell_{p}(\xi)$
which however may not be continuous across the elements. This is illustrated in Figure 1b.
@@
Modify the flux approximation $\bm{F}^{\ast\delta}_{h}(\xi)$ so that it becomes continuous across the elements. Let $\bm{F}^{\ast}_{{e+\frac{1}{2}}}$ be some numerical flux function that approximates the flux $\bm{F}^{\ast}$ at $x=x_{{e+\frac{1}{2}}}$. Then the continuous flux approximation is given by
$\bm{F}^{\ast}_{h}(\xi)=\left[\bm{F}^{\ast}_{{e-\frac{1}{2}}}-\bm{F}^{\ast\delta}_{h}(0)\right]g_{L}(\xi)+\bm{F}^{\ast\delta}_{h}(\xi)+\left[\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{F}^{\ast\delta}_{h}(1)\right]g_{R}(\xi)$
@@
$\mathsf{\partial_{\xi}\bm{F}^{\ast}_{h}}=\left[\bm{F}^{\ast}_{e-\frac{1}{2}}-\mathsf{V}_{L}^{\top}\mathsf{\bm{F}^{\ast}}\right]\mathsf{b}_{L}+\mathsf{D}\mathsf{\bm{F}^{\ast}}+\left[\bm{F}^{\ast}_{e+\frac{1}{2}}-\mathsf{V}_{R}^{\top}\mathsf{\bm{F}^{\ast}}\right]\mathsf{b}_{R},\qquad\mathsf{b}_{L}=\begin{bmatrix}g_{L}^{\prime}(\xi_{0})\\ \vdots\\ g_{L}^{\prime}(\xi_{N})\end{bmatrix},\qquad\mathsf{b}_{R}=\begin{bmatrix}g_{R}^{\prime}(\xi_{0})\\ \vdots\\ g_{R}^{\prime}(\xi_{N})\end{bmatrix}$
@@
$\mathsf{\partial_{\xi}\bm{F}^{\ast}_{h}}=\bm{F}^{\ast}_{{e-\frac{1}{2}}}\mathsf{b}_{L}+\mathsf{D}_{1}\mathsf{\bm{F}^{\ast}}+\bm{F}^{\ast}_{{e+\frac{1}{2}}}\mathsf{b}_{R},\qquad\mathsf{D}_{1}=\mathsf{D}-\mathsf{b}_{L}\mathsf{V}_{L}^{\top}-\mathsf{b}_{R}\mathsf{V}_{R}^{\top}$ (11)